Two automobiles are 150 kilometer apart and traveling toword each other.one automoblie is moving at 60km/h and the other is moving at 40km/h mph.in how many hours will they meet?
1
Expert's answer
2018-04-19T09:47:08-0400
Time to meet of automobiles
t=l/(v_1+v_2 )
t=150/(60+40)=1.5 hours